OpenStudy (anonymous):

What is the simplified form of (2x/3x+9) / (5x/x^2-x-12)?

OpenStudy (anonymous):

@sami-21 help ?! :)

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

\[\huge \frac{\frac{2x}{3x + 9}}{\frac{5x}{x^2 - x - 12}} \implies \frac{2x \cdot (x^2 - x - 12)}{5x \cdot (3x + 9)}\] Factorize numerator first..

OpenStudy (anonymous):

Can you factorize it ? \[x^2 - x - 12\]

OpenStudy (anonymous):

(x-4)(x+3)

OpenStudy (anonymous):

Great..

OpenStudy (anonymous):

Now what

OpenStudy (anonymous):

From denominator you can factor out 3..

OpenStudy (anonymous):

By use of distributive property: Factor out 3 from it: \[3x + 9\]

OpenStudy (anonymous):

I got 2x (x-4) / 15x

OpenStudy (anonymous):

Yes.. You can cancel x also here..

OpenStudy (anonymous):

\[\frac{2x(x-4)}{15x} \implies \frac{2}{15}(x-4)\]

OpenStudy (anonymous):

Getting?? @jrosesweet

OpenStudy (anonymous):

Yes so that's the answer?

OpenStudy (anonymous):

Yes that is the answer.. Or you can write it as : \[\frac{2x - 8}{15}\] If you want..

OpenStudy (anonymous):

Thanks!! @waterineyes